# Fayet–Iliopoulos IR infrared

### v2

@@
low-energy effective field theory
@@ 1 Introduction
Far from being a residual geometric artifact, modular symmetry plays an active role in shaping effective theories from string compactifications. It constrains the structure of Yukawa couplings, governs the form of non-perturbative superpotentials, and serves as a guiding principle for constructing and validating complete low-energy effective field theorys. [^1] Modular invariance was explicitly used to construct the full non-linear supergravity LEEFT in concrete heterotic orbifold models in [7]; a similar principle guided the derivation of anti-D3 uplifted flux compactifications in [8].
@@ 2.1.4 Supersymmetric stabilisation of matter fields
To retain supersymmetry at or close to the compactification scale, one must demand $\langle{D_{\text{anom}}}\rangle=0$, which leads to non-vanishing VEVs, $\langle{\Phi_{\alpha}^{(n_{\alpha})}}\rangle$. Furthermore, it is known that also the $F$-terms can be simultaneously cancelled with $D$-flat VEV configurations (while also $\langle{W}\rangle\sim 0$) [98].
@@ 2.2 Modular-Invariant Effective Action
$S\penalty 10000\ \xmapsto{\penalty 10000\ \penalty 10000\ }\penalty 10000\ S-\frac{\delta_{\text{GS}}}{8\pi^{2}}\ln(\mathrm{i}c\tau+d)\,$, (21)
@@
$K\penalty 10000\ \xmapsto{\penalty 10000\ \gamma\penalty 10000\ }\penalty 10000\ K+\log|\mathrm{i}c\tau+d|^{2}\,,\qquad W\penalty 10000\ \xmapsto{\penalty 10000\ \gamma\penalty 10000\ }\penalty 10000\ \frac{W}{(\mathrm{i}c\tau+d)}\,$, (22)
@@ 2.2.2 The Superpotential
with $\eta(\tau)$ the Dedekind eta function defined in Equation 16 and $H_{a}(\tau)$ modular invariant functions defined as
$H_{a}(\tau)=\left(\frac{E_{4}(\tau)}{\eta^{8}(\tau)}\right)^{n_{a}}\left(\frac{E_{6}(\tau)}{\eta^{12}(\tau)}\right)^{m_{a}}P(j(\tau))\penalty 10000\ =\penalty 10000\ (j(\tau)-1728)^{m_{a}/2}(j(\tau)^{n_{a}/3}P\left(j(\tau)\right)\,$, (29)
where $n_{a},\,m_{a}$ are some non-negative integers, $j(\tau)$ is the Klein invariant function given by
@@ 3 Mapping the modular de Sitter landscape
We assume that the matter fields acquire VEVs of the same order, taken to be $\mathcal{O}(10^{-1})$ so as to remain within the regime of validity of the Kähler potential calculation. Such VEVs can lead to the parameter values $\tilde{\lambda}_{1}=3/25000$ and $\tilde{\lambda}_{2}=3/5000000$ in Equation 25. We also set $c_{1}=1$ and $c_{2}=8\pi^{2}\mathrm{e}$, so that the coefficients in Equation 32 differ in magnitude.
The parameters $b_{1}$, $b_{1}^{0}$, and $\delta_{\text{GS}}$ are computed from the explicit orbifold model described in Appendix A for one $\mathrm{SU}(3)$ factor, while $b_{2}$ and $b_{2}^{0}$ are chosen to ensure that the second condensate becomes dominant at a different order in moduli expansion. Since the Kač-Moody level is $k=1$ for standard heterotic orbifold models, we take $k_{1}=k_{2}=1$. The decoupling scale $M_{d}$ must lie not far below the Planck scale; [^9] As shown in [106], exotic matter can decouple a few orders of magnitude below the Planck scale. here we fix it to be $M_{d}=1/65$. Given these choices, the detailed structure of the critical points is controlled by the integers $m_{a}$ and $n_{a}$ in the modular invariant functions (29). All parameter values used in our analysis are summarised in Table 2.
@@
As mentioned earlier, instead of focusing only on local minima, our search aims to locate as many critical points of the scalar potential in Equation 36 as possible. For each sampled point, we compute $\nabla V$ and use FindRoot to solve for critical configurations. If $\nabla V$ fails to vanish within tolerance (or any numerical issues arise during evaluation), the point is classified as Not a critical point.
@@
Further, the resulting distributions of the various classes are shown in the complex planes of $\tau$ and $S$ in Figures 5 and 6, respectively, where the marker shape and colour denote the class, and opacity is adjusted to highlight vacuum and tachyonic gauge coupling classes. A striking feature emerges in Figure 5: solutions in the Tachyonic axion class cluster in angular regions forming a distinctive “penacho” [^11] The name “penacho” is a nod to the famed feathered headdress of Moctezuma, currently held in Vienna; it is a cultural reminder that what is displaced can still speak meaningfully to its origins. pattern in the $\tau$-plane. The significance of this plume-like structure remains unclear, though it may suggest deeper constraints imposed by modular invariance. Further investigation is needed to determine its potential relevance. Stable vacua appear more sparsely, typically near the boundaries of the fundamental domain. While Not a critical point outcomes appear to occupy complementary regions, this effect is not statistically significant.
@@ 3.1.2 Unstable dS
To conclude this subsection, we investigate whether modular symmetry plays an essential role in our results. To this end, we explicitly break modular invariance in the potential and repeat our search. We find that the appearance of tachyonic axion saddle points is strongly suppressed: out of $10^{5}$ candidate critical points, only eight such saddles occur without modular invariance, compared to $970$ in the modular-invariant case. This striking difference suggests that modular symmetry is not merely a consistency requirement, but actively shapes the vacuum structure and the presence of cosmologically relevant saddles.
@@ 3.2.1 Refined dS conjecture
where $c$ and $c^{\prime}$ are positive constants of $\mathcal{O}(1)$, and $\mathrm{min}(\nabla^{i}\nabla_{j}V)$ denotes the smallest eigenvalue of the Hessian in an orthonormal frame.
@@
to verify whether it meets the required threshold. Our analysis shows that all unstable dS vacua in our model satisfy this bound, with values reaching as low as $r\lesssim-1.9$.
@@ 4 Multifield quintessence from modular dS saddles
Motivated by this, we now investigate cosmological evolution in the modular invariant heterotic potentials introduced above. In particular, we study the dynamics near a dS saddle point in the full four-field system of the complex dilaton $S$ and complex structure modulus $\tau$. This extends earlier single-field hilltop constructions to a setting where multiple moduli can roll simultaneously, as is generically expected in string compactifications.
@@
$\psi^{\tau}$ $9.155\times 10^{-117}$ $9.468\times 10^{-117}$
$\psi^{S}$ $7.528\times 10^{-120}$ $9.527\times 10^{-120}$
@@
Table 3: Axionic saddle point of the potential (36). The mass eigenstate $\psi^{I}$ has a dominant contribution of the modulus $\Psi^{I}$. The unstable direction is mostly governed by the axionic fields $\operatorname{Re}(\tau)$ and $\operatorname{Im}(S)$. All VEVs and masses are given in reduced Planck units with $M_{\mathrm{Pl}}=1$.
@@ 4.2 The fate of dark energy: descent into adS
$\psi^{\tau}$ $5.023\times 10^{-117}$ $5.009\times 10^{-117}$
$\psi^{S}$ $1.023\times 10^{-118}$ $1.006\times 10^{-118}$

### v3

# arXiv:2509.22781v3[hep-th] 01 Mar 2026 BU bottom-up

@@
low-energy effective field theorie
@@ 1 Introduction
Far from being a residual geometric artifact, modular symmetry plays an active role in shaping effective theories from string compactifications. It constrains the structure of Yukawa couplings, governs the form of non-perturbative superpotentials, and serves as a guiding principle for constructing and validating complete low-energy effective field theories. [^1] Modular invariance was explicitly used to construct the full non-linear supergravity LEEFT in concrete heterotic orbifold models in [7]; a similar principle guided the derivation of anti-D3 uplifted flux compactifications in [8].
@@ 2.1.4 Supersymmetric stabilisation of matter fields
To retain supersymmetry at or close to the compactification scale, one must demand $\langle{D_{\text{anom}}}\rangle=0$, which leads to non-vanishing VEVs, $\langle{\Phi_{\alpha}^{(n_{\alpha})}}\rangle$. Furthermore, it is known that also the $F$-terms can be simultaneously cancelled with $D$-flat VEV configurations (while also $\langle{W}\rangle\sim 0$) [98, 99, 100]. Due to the large amount of fields solving these equations, there is enough freedom to fix by hand some of the VEVs without significantly affecting the results. However, it could be interesting to study this in detail, which will be done in an extension of this work.
@@ 2.2 Modular-Invariant Effective Action
$S\penalty 10000\ \xmapsto{\penalty 10000\ \penalty 10000\ }\penalty 10000\ S-\frac{\delta_{\text{GS}}}{8\pi^{2}}\log(c\tau+d)\,$, (21)
@@
$K\penalty 10000\ \xmapsto{\penalty 10000\ \gamma\penalty 10000\ }\penalty 10000\ K+\log|c\tau+d|^{2}\,,\qquad W\penalty 10000\ \xmapsto{\penalty 10000\ \gamma\penalty 10000\ }\penalty 10000\ \frac{W}{(c\tau+d)}\,$, (22)
@@ 2.2.2 The Superpotential
with $\eta(\tau)$ the Dedekind eta function defined in Equation 16 and $H_{a}(\tau)$ are the most general modular invariant functions without singularities in the fundamental domain, parameterised by [108, 109]
$H_{a}(\tau)=\left(\frac{E_{4}(\tau)}{\eta^{8}(\tau)}\right)^{n_{a}}\left(\frac{E_{6}(\tau)}{\eta^{12}(\tau)}\right)^{m_{a}}P(j(\tau))\penalty 10000\ =\penalty 10000\ (j(\tau)-1728)^{m_{a}/2}j(\tau)^{n_{a}/3}P\left(j(\tau)\right)\,$, (29)
where $n_{a},\,m_{a}$ are some non-negative integers, $P$ is a polynomial function of the Klein invariant function $j(\tau)$ given by
@@ 3 Mapping the modular de Sitter landscape
We assume that the matter fields acquire VEVs of the same order, taken to be $\mathcal{O}(10^{-1})$ so as to remain within the regime of validity of the Kähler potential calculation. Such VEVs can lead to the parameter values $\tilde{\lambda}_{1}=3/25000$ and $\tilde{\lambda}_{2}=3/5000000$ in Equation 25. We also set $c_{1}=1$ and $c_{2}=8\pi^{2}\mathrm{e}$, so that the coefficients in Equation 32 differ in magnitude. [^9] These values represent a convenient choice and are not unique; however, we have explicitly verified that reasonable variations of $c_{1}$ and $c_{2}$ do not modify the qualitative features of our results.
The parameters $b_{1}$, $b_{1}^{0}$, and $\delta_{\text{GS}}$ are computed from the explicit orbifold model described in Appendix A for one $\mathrm{SU}(3)$ factor, while $b_{2}$ and $b_{2}^{0}$ are chosen to ensure that the second condensate becomes dominant at a different order in moduli expansion. Since the Kač-Moody level is $k=1$ for standard heterotic orbifold models, we take $k_{1}=k_{2}=1$. The decoupling scale $M_{d}$ must lie not far below the Planck scale; [^10] As shown in [98], exotic matter can decouple a few orders of magnitude below the Planck scale. here we fix it to be $M_{d}=1/65$. Given these choices, the detailed structure of the critical points is controlled by the integers $m_{a}$, $n_{a}$ and by the polynomial $P(j(\tau))$ in the modular invariant functions (29), for computational simplicity, the polynomial was fixed [^11] Other choices may modify the shape of the potential, leading to other results. However, studying all the possible choices for $P(j(\tau))$ is beyond the scope of this work. as $P(j(\tau))=1$ and $m_{a}$, $n_{a}$ were taken [^12] Since $H(\tau)$ is a holomorphic singularity-free function in the fundamental domain, it admits a locally convergent series expansion. Hence, taking higher values for $n_{a}$ and $m_{a}$ yields contributions that are negligible compared to those from smaller values. within $[0,12]$. All parameter values used in our analysis are summarised in Table 2.
@@
As mentioned earlier, instead of focusing only on local minima, our search aims to locate as many critical points of the scalar potential in Equation 36 as possible. For each sampled point, we compute $\nabla V$ and use FindRoot to solve for critical configurations. Numerical convergence was established by requiring $|\nabla V|\lesssim 10^{-20}$ for all critical points. If $\nabla V$ fails to vanish within tolerance (or any numerical issues arise during evaluation), the point is classified as Not a critical point.
@@
Further, the resulting distributions of the various classes are shown in the complex planes of $\tau$ and $S$ in Figures 5 and 6, respectively, where the marker shape and colour denote the class, and opacity is adjusted to highlight vacuum and tachyonic gauge coupling classes. Concerning the possible presence of modularly equivalent points in the numerical results, we carried out an explicit verification for modular redundancies. While search seeds were drawn from the fundamental domain, we performed an a posteriori verification to ensure the final dataset is free of modular redundancies. Equivalent solutions identified via the $\mathrm{SL}(2,\mathds{Z})$ mapping (including the Green-Schwarz transformation) constituted a negligible subset ($0.01\%$) and were subsequently removed. A striking feature emerges in Figure 5: solutions in the Tachyonic axion class cluster in angular regions forming a distinctive “penacho” [^14] The name “penacho” is a nod to the famed feathered headdress of Moctezuma, currently held in Vienna; it is a cultural reminder that what is displaced can still speak meaningfully to its origins. pattern in the $\tau$-plane. The significance of this plume-like structure remains unclear, though it may suggest deeper constraints imposed by modular invariance. Further investigation is needed to determine its potential relevance. Stable vacua appear more sparsely, typically near the boundaries of the fundamental domain. While Not a critical point outcomes appear to occupy complementary regions, this effect is not statistically significant.
@@ 3.1.2 Unstable dS
To conclude this subsection, we investigate whether modular symmetry plays an essential role in our results. To this end, we explicitly break modular invariance in the potential [^16] To this end, we modified the non-perturbative contribution to the superpotential (28) as $W_{\mathrm{gc}}=\Omega_{1}(S)H_{1}(\tau)+\Omega_{2}(S)H_{2}(\tau)$. In this way, the superpotential (35) does not transform as a modular form with modular weight 1-1, breaking modular invariance. and repeat our search. We find that the appearance of tachyonic axion saddle points is strongly suppressed: out of $10^{5}$ candidate critical points, only eight such saddles occur without modular invariance, compared to $970$ in the modular-invariant case. This striking difference suggests that modular symmetry is not merely a consistency requirement, but actively shapes the vacuum structure and the presence of cosmologically relevant saddles.
@@ 3.2.1 Refined dS conjecture
where $c$ and $c^{\prime}$ are positive constants of $\mathcal{O}(1)$, and $\mathrm{min}(\nabla^{i}\nabla_{j}V)$ denotes the smallest eigenvalue of the field-space covariant mass matrix $\nabla^{i}\nabla_{j}V$.
@@
to verify whether it meets the required threshold. Our analysis shows that all unstable dS vacua in our model satisfy this bound, with values reaching as low as $r\lesssim-11$.
@@ 4 Multifield quintessence from modular dS saddles
Motivated by this, we now investigate cosmological evolution in the modular invariant heterotic potentials introduced above. In particular, we study the dynamics near a dS saddle point in the full four-field system of the complex dilaton $S$ and Kähler modulus $\tau$. This extends earlier single-field hilltop constructions to a setting where multiple moduli can roll simultaneously, as is generically expected in string compactifications.
@@
$\psi^{\tau}$ $3.190\times 10^{-116}$ $3.292\times 10^{-116}$
$\psi^{S}$ $2.830\times 10^{-119}$ $3.574\times 10^{-119}$
@@
Table 3: Axionic saddle point of the potential (36). The values of $m^{2}$ correspond to the eigenvalues of the mass matrix $\nabla^{i}\nabla_{j}V$. The mass eigenstate $\psi^{I}$ has a dominant contribution of the modulus $\Psi^{I}$. The unstable direction is mostly governed by the axionic fields $\operatorname{Re}(\tau)$ and $\operatorname{Im}(S)$. All VEVs and masses are given in reduced Planck units with $M_{\mathrm{Pl}}=1$.
@@ 4.2 The fate of dark energy: descent into adS
$\psi^{\tau}$ $1.925\times 10^{-116}$ $1.920\times 10^{-116}$
$\psi^{S}$ $5.292\times 10^{-118}$ $5.206\times 10^{-118}$
@@ Appendix A A concrete orbifold model
At leading order, this approximation includes $11$ out of a set of $\mathcal{O}(100)$ matter fields. All additional states appear at higher orders and are hence further suppressed. This implies that their contributions to $W_{\text{Yuk}}$ can be ignored at first order.